### v1

# CallNavi: A Study and Challenge on Function Calling Routing and Invocation in Large Language Models

Abstract
Interacting with a software system via a chatbot can be challenging, especially when the chatbot needs to generate API calls, in the right order and with the right parameters, to communicate with the system. API calling in chatbot systems poses significant challenges, particularly in complex, multi-step tasks requiring accurate API selection and execution. We contribute to this domain in three ways: first, by introducing a novel dataset designed to assess models on API function selection, parameter generation, and nested API calls; second, by benchmarking state-of-the-art language models across varying levels of complexity to evaluate their performance in API function generation and parameter accuracy; and third, by proposing an enhanced API routing method that combines general-purpose large language models for API selection with fine-tuned models for parameter generation and some prompt engineering approach. These approaches lead to substantial improvements in handling complex API tasks, offering practical advancements for real-world API-driven chatbot systems.
CallNavi: A Study and Challenge on Function Calling Routing and Invocation in Large Language Models
Yewei Song1, Cedric Lothritz2, Xunzhu Tang1, Saad Ezzini3, Jacques Klein1, Tegawendé F. Bissyandé1, Andrey Boytsov4, Ulrick Ble 4, Anne Goujon4, 1University of Luxembourg, 2Luxembourg Institute of Science and Technology, 3Lancaster University, 4BGL BNP PARIBAS, Correspondence: yewei.song@uni.lu
1 Introduction
Modern conversational AI systems, such as chatbots, rely heavily on accurate API calling to enable effective user interactions. In complex scenarios, models must go beyond generating simple API calls to handle challenges such as selecting the correct API from an extensive list, orchestrating multiple sequential API calls, and managing nested calls. Despite significant progress in generating syntactically correct single API calls, there has been limited focus on tasks requiring a sequence of API calls with logical dependencies, a crucial capability for real-world applications.
Large Language Models (LLMs) such as GPT-4 Achiam et al. (2023) and Llama Touvron et al. (2023) have demonstrated remarkable capabilities across a wide range of natural language processing tasks. By leveraging extensive pretraining on diverse datasets, LLMs can generate coherent and contextually relevant responses, making them essential components of modern conversational AI systems. However, while their abilities in tasks like summarization and question answering are well-established, their proficiency in generating structured outputs—such as API calls, program code, and other machine-readable formats—remains a challenging frontier.
Structured output generation requires LLMs to produce outputs that adhere to predefined syntactic and semantic rules, a task significantly more constrained than generating free-form text. Recent works have explored structured output generation for applications like code generation Chen et al. (2021), table completion Herzig et al. (2020), and multi-turn dialogue Budzianowski et al. (2018). Tools like CodeX Chen et al. (2021) and AlphaCode Li et al. (2022) have specifically focused on structured programming tasks, demonstrating the potential of LLMs in generating functionally valid code. Similarly, methods such as chain-of-thought (CoT) prompting Wei et al. (2022) and tool-augmented reasoning frameworks Qin et al. (2023) have enabled LLMs to reason step by step, improving their accuracy in complex tasks.
Several earlier works have explored function calling and tool-use capabilities in LLMs, laying a foundation for modern benchmarks. Early studies, such as BotBase Zamanirad et al. (2017), focused on translating natural language into API calls, demonstrating initial progress in automating tool use. More recent benchmarks, including API-Bank Li et al. (2023), ToolEyes Ye et al. (2024), and BFCL Yan et al. (2024), evaluate LLMs on their ability to execute API calls. For instance, API-Bank assesses tool-augmented models on real-world tasks but operates with fewer than five API candidates per task, limiting scenario complexity. Similarly, ToolBench Liu et al. (2024) and ToolEyes test multi-tool scenarios but lack support for nested or highly interdependent API calls.
To address these gaps, we propose CallNavi, a novel benchmark designed to evaluate large language models (LLMs) on tasks involving:
  • API selection from an unfiltered list of over 100 candidates;
  • Execution of multiple sequential API calls;
  • Handling of nested API interactions.
Unlike prior datasets, CallNavi introduces real-world complexity by simulating unfiltered API selection and combining generation with routing tasks. It categorizes questions into easy, medium, and hard levels, enabling the evaluation of models across varying levels of task complexity. Our benchmark also proposes new metrics, including stability score, to measure consistency in model predictions over multiple runs.
We comprehensively benchmark 17 LLMs on CallNavi, covering diverse commercial, general-purpose, and fine-tuned models. Our findings reveal key insights into the strengths and limitations of current models, providing a foundation for further advancements in API selection and function calling.
2 Related Work
2.1 Function Calling and API Selection
Table 1 compares CallNavi with existing benchmarks across key dimensions, highlighting its unique contributions. Notably, our CallNavi is the only benchmark to test models on unfiltered API selection with over 100 candidates. And CallNavi incorporate both multi-call and nested API tasks in a single dataset, provide a diverse set of 10 domains while maintaining structured difficulty levels.
Table 1: Comparison of CallNavi with existing API function-calling benchmarks test set.
Benchmark Domains Questions Max API Multi-Call Nested
Candidates
CallNavi 10 729 115 Yes Yes
API-Bank 8 753 $<$5 Yes Yes
ToolEyes 41 382 $<$20 Yes No
ToolBench 8 795 32 Yes No
BFCL (API) N/A 70 $<$5 Yes No
2.2 Techniques for Structured Output and Stability
Recent advances in structured output generation, such as MetaGPT and CodeAgent, have inspired innovations in generating API call sequences Hong et al. (2023); Tang et al. (2024). These works emphasize task decomposition and step-by-step reasoning, which align with our use of function-calling routing for complex API tasks.
Ensuring stability across multiple runs remains a critical challenge for function-calling tasks. Prior works addressing LLM stability Du et al. (2023); Tian et al. (2023) have informed our development of the stability score, which quantifies consistency in predictions. This metric complements traditional accuracy metrics like AST match and exact match, enabling a more comprehensive evaluation of model reliability in real-world scenarios.
Figure 1: Pipeline of API calling while using LLMs
3 Research Questions
  • Benchmark Which LLMs have the best performance for function calling in a real-world scenario?
  • Evaluation Which is the best way to evaluate the API function calling ability of LLMs?
  • Optimization How to enhance API function calling ability for zero/few-shot LLM?
4 CallNavi Dataset
To create our dataset, we adopted a hybrid approach that combines automated generation with manual validation and construction to ensure high-quality and diverse data across different levels of difficulty. The process consisted of the following steps:
Initial API Function Generation.
Using GPT-4o, we generated API function names, descriptions, parameters, and return values based on a variety of scenario descriptions spanning multiple domains. These domains were selected to reflect realistic use cases across 10 common chatbot application areas, as described in Table 2. This ensures that the dataset evaluates CallNavi in scenarios requiring advanced task routing, contextualization, and regulatory compliance.
Validation and Refinement.
All generated API functions were manually reviewed for accuracy, consistency, and relevance. Specifically:
  • Parameters were checked to ensure they aligned with real-world API design conventions.
  • Ambiguities or redundancies in function descriptions were resolved.
  • Naming conventions for parameters and return values were standardized to ensure consistency across the dataset.
Generation of Easy Questions.
For the easy subset, we used GPT-4o to generate questions related to API usage. These questions were subsequently validated to ensure:
  • Relevance to the provided APIs,
  • Syntactic and semantic correctness, and
  • Coverage of straightforward, single-function usage scenarios.
Manual Construction of Medium and Hard Questions.
Medium and hard questions were manually crafted to reflect increasingly complex API calling scenarios. The criteria and considerations for these levels were as follows:
  • Medium Questions: Focused on multi-step tasks requiring the use of multiple APIs in sequence. These tasks test the model’s ability to identify dependencies between API calls while maintaining logical flow.
  • Hard Questions: Designed to address edge cases, ambiguous queries, and nested API calls requiring advanced reasoning or parameter transformations. Scenarios simulate real-world challenges, such as incomplete user inputs or conflicting requirements.
Examples of complex scenarios from medium and hard questions are included in Appendix E.
Quality Control.
The dataset underwent a multi-stage quality assurance process to ensure its reliability:
  • Each generated instance was cross-checked by multiple annotators for correctness and consistency.
  • For manually written instances, authors verified adherence to the design criteria.
  • Errors, ambiguities, and inconsistencies were flagged and resolved iteratively.
Summary.
The CallNavi dataset combines automation with human oversight, resulting in a benchmark that is both realistic and challenging. By spanning easy, medium, and hard tasks across diverse real-world domains, as outlined in Table 2, the dataset evaluates LLM capabilities in scenarios requiring robust task routing, contextual understanding, and API management. The dataset format and metadata details are provided in Appendix A.
Table 2: CallNavi dataset domains, questions and difficulties statistics table.
{tblr}
cell11 = r=2, cell12 = r=2, cell13 = r=2, cell14 = c=3c, vlines, hline1,3-14 = -, hline2 = 4-6, Domain & API
Functions Questions Difficulty
Easy Medium Hard
Banking 91 115 70 28 17
Shopping 81 65 41 17 7
Logistics 46 65 40 17 8
Aviation 48 80 44 24 12
Healthcare 20 47 31 10 6
Public Services 82 85 50 27 8
Human Resources 20 35 21 13 1
Hotel Industry 49 65 40 19 6
Insurance 42 60 40 11 9
Telecommunications 100 112 79 22 11
Overall 579 729 456 188 85
4.1 Dataset
The CallNavi dataset [^1] https://anonymous.4open.science/r/CallNavi-783E evaluates LLMs’ task routing and API calling capabilities across multiple domains. As shown in Table 2, it contains 729 questions of varying difficulty and API interaction complexity, along with 579 distinct API functions. Questions are categorized into easy, medium, and hard levels:
  • Easy (456 questions): Require a single API call to fulfill the task.
    Example: A user checking their bank account balance with one straightforward API call.
  • Medium (188 questions): Involve multiple APIs within the same question, with all necessary parameters provided in the context.
    Example: A shopping query needing product details and stock availability via two independent API calls.
  • Hard (85 questions): Require multiple API calls where some parameters depend on responses from previous calls, adding complexity. 5 steps maximum of API.
    Example: Updating delivery status by first retrieving a package ID, then using it to fetch the delivery status through sequential API calls.
This dataset tests LLMs’ ability to perform function-calling routing and parameters generation across varying difficulties, assessing both basic single-call handling and complex multi-step nested requests. Zero-shot and few-shot models must infer correct API interactions only from the question context from different difficulties.
5 Metrics and Evaluation
5.1 API Parameters AST Match
Our study utilizes Abstract Syntax Tree (AST) evaluation to assess models’ ability to generate accurate JSON outputs for API calls. The format of the output JSON and parameters follows a structure similar to the BFCL dataset in Section A Yan et al. (2024). We parse the generated JSON string into an object and compare each component with the ground truth, such as the API list and parameters.
In scenarios involving multi-step API calls where parameters depend on previous steps or where text inputs may not have a single definitive answer, placeholder tokens are used for parameters. These tokens are positionally aligned with the ground truth, and we exclude them from strict comparisons during evaluation.
Our AST evaluation process is based on three key criteria, examples in Figure 2:
  • Syntax Validity: Whether the JSON string can be correctly parsed into an object without syntax errors.
  • Structural Accuracy: Whether the parsed API calls match the ground truth and include the correct parameter names(keys).
  • AST Exact Match: Whether the entire parsed object, including its structure and content(values), is identical to the ground truth.
Figure 2: Example of evaluation pipeline
As Figure 3 shows, we begin by checking the syntax validity of the generated JSON structure. If syntax errors are detected, we apply a JSON fix prompt to repair the structure or convert alternative formats (such as function-calling syntax) into valid JSON. Once the structure is valid, we assess structural accuracy by comparing the predicted JSON with the ground truth. A structural match is scored as 1. Finally, we convert both the predicted and ground truth JSONs into object trees, comparing each node and leaf. A perfect match across all nodes results in a score of 1 for AST Exact Match.
This multi-step evaluation ensures a thorough assessment of both the accuracy of API function calls and the structural integrity of the parameters, allowing for a granular analysis of the model’s performance.
Figure 3: Pipeline of AST Match Score
5.2 LLM-as-a-Judge Evaluation
We also use GPT-4o language models to evaluate whether the generated JSON outputs correspond accurately to the ground truth Zheng et al. (2023). This approach aims to observe if LLMs can perform such evaluation tasks with high precision. By leveraging an LLM for this purpose, we assess its capability to compare and validate structured data, thereby determining its effectiveness in automating the evaluation process. An example prompt is shown in appendix B.2.
5.3 Stability Score
In chatbot systems, consistent outputs for identical inputs are crucial to ensure reliability and user trust, especially in professional settings. Users expect the same accurate response each time they ask the same question. Inconsistencies can cause confusion, erode confidence, and lead to errors, particularly in critical fields like finance or healthcare.
We propose an Election Stability Score to evaluate the consistency of API outputs across multiple runs for the same input. This score mirrors an election process, selecting the majority output as the final answer. To calculate the score, we define:
  • $N$: Total number of outputs (samples).
  • $F_{1}$: Maximum frequency among all unique outputs (count of the most frequent output).
  • $F_{2}$: Second maximum frequency (count of the second most frequent output).
The stability score is calculated as:
Stability $\text{Stability Score}=\frac{F_{1}-F_{2}}{N-F_{2}}$
This quantifies the consistency of the model’s outputs: If there’s a tie for the most frequent output ($F_{1}=F_{2}$), the stability score is set to 00, indicating no consensus; If the most frequent output is unique ($F_{1}>F_{2}$), the score ranges from 00 to $1$, reflecting the dominance of the most frequent output.
To ensure reliable comparisons and reduce errors, we preprocess the outputs by removing unnecessary spaces, newlines, and formatting inconsistencies, converting the text to lowercase, and stripping extraneous characters that could cause mismatches. More detailed analysis can be found in the appendix D.
While ‘freq@topk‘ is often used to evaluate model performance, it fails to capture stability in the generated outputs. For example, if a model produces the sequence "AABBC" across multiple runs, ‘freq@topk‘ might assign a high score of 0.4 because the most frequent token ("A") appears 40% of the time. However, this sequence is not stable, as no single output consistently dominates. In contrast, our stability score focuses on the dominance of the most frequent output, offering a better measure of a model’s reliability in structured tasks.
To further evaluate the stability of the model’s outputs, we calculate the average Levenshtein distance between the first answer and each subsequent outputZhang et al. (2017). We normalize Levenshtein distance using the following formula.
$Score_{Lev}=\frac{1}{n}\sum_{i=1}^{n}(1-\frac{lev(x_{0},x_{i})}{Max(len(x_{0}),len(x_{i}))})$
6 Experiments and Benchmark
6.1 Models
To evaluate the benchmark, we selected models based on their performance, architecture, and relevance to function-calling tasks. The selection criteria focused on general-purpose and fine-tuned models optimized for function calling or JSON generation, ensuring a well-rounded comparison between zero-shot and fine-tuned capabilities. Models like BART and traditional retrieval-based approaches were excluded as they lack the ability to select APIs from extensive lists or structure API calls sequentially, which are critical for the complexity of this task.
Table 3 organizes the selected models into four groups: commercial models, middle-large models (10B+ parameters), small models (5B–10B parameters), and lightweight models (under 5B parameters).
Model Name Origin Parameters
GPT-4o OpenAI Achiam et al. (2023) N/A
GPT-4o-mini OpenAI Achiam et al. (2023) N/A
Gemini 1.5 Flash Google Reid et al. (2024) N/A
LLaMA 3.1 Meta AI Meta-AI (2024a) 70B
Command-R Command AI Cohere 35B
Gemma2 Google Team et al. (2024) 27B
Mistral-Small Mistral AI Ai (2024a) 22B
Phi3 Microsoft Abdin et al. (2024) 14B
Mistral-Nemo Mistral AI Ai (2024b) 12B
Gemma2 Google Team et al. (2024) 9B
LLaMA 3.1 Meta AI Meta-AI (2024a) 8B
xLAM Salesforce Zhang et al. (2024) 7B
NemoTron-Mini NVIDIA Paul (2024) 4B
Phi3.5 Microsoft Trufinescu (2024) 3B
LLaMA 3.2 Meta AI Meta-AI (2024b) 3B
NexusRaven Nexusflow.ai Nexusflow.ai (2023) 13B
Gorilla Berkeley Patil et al. (2023) 7B
Table 3: Comparison of General Purpose LLMs.
We choose 2 fine-tuned Function Calling models for testing which have top performance on the BFCL leaderboard: NexusRaven and Gorilla OpenFunctions v2. Then we found that Firefunction v2 on the leaderboard cannot accept that long API calling list in the pilot test Garbacki and Chen (2024).
{tblr}
row1 = c, row2 = c, cell11 = r=2, cell12 = r=2, cell13 = c=4, cell17 = r=2, cell18 = r=2, cell19 = c=5, cell114 = r=2, cell31 = r=3, cell61 = r=6, cell121 = r=3, cell151 = r=3, cell181 = r=2, cell201 = r=3, hline1,23 = -0.08em, hline2 = 3-6,9-13, hline3 = -, Category & Models API Calling Routing
Exact Match Syntax
Validity Structural
Accuracy API Calling with Parameters
AST Match Overall
GPT Score
Easy Medium Hard All Easy Medium Hard All Avg. Macro Avg.
Commercial
Models GPT4o 0.978 0.914 0.611 0.919 0.993 0.887 0.802 0.638 0.388 0.711 0.609 0.913
GPT4o mini 0.971 0.930 0.564 0.913 0.994 0.869 0.800 0.648 0.364 0.710 0.604 0.908
Gemini 1.5 Flash 0.973 0.904 0.564 0.908 0.945 0.806 0.728 0.462 0.258 0.604 0.483 0.876
Large
General
LLMs LLAMA3.1 70B 0.945 0.835 0.470 0.861 0.967 0.299 0.296 0.191 0.094 0.245 0.194 0.583
CommandR 35B 0.789 0.877 0.529 0.781 0.969 0.189 0.167 0.095 0.047 0.134 0.103 0.400
Gemma2 27B 0.945 0.877 0.552 0.882 0.982 0.226 0.217 0.143 0.070 0.181 0.143 0.476
Mistral-Small 22B 0.885 0.819 0.494 0.823 0.986 0.196 0.201 0.106 0.059 0.160 0.122 0.417
Phi3 14B 0.050 0.032 0.011 0.041 0.283 0.021 0.019 0.010 0.0 0.015 0.010 0.082
Mistral-Nemo 12B 0.927 0.808 0.470 0.843 0.842 0.271 0.296 0.127 0.035 0.222 0.153 0.524
Middle
LLMs Gemma2 9B 0.962 0.845 0.506 0.879 0.983 0.220 0.241 0.095 0.059 0.182 0.132 0.488
LLAMA3.1 8B 0.916 0.813 0.552 0.847 0.925 0.207 0.223 0.058 0.059 0.162 0.113 0.422
xLAM-fc 7B 0.642 0.377 0.188 0.521 0.990 0.271 0.307 0.117 0.058 0.229 0.161 0.554
Light
Models nemotron-mini 4B 0.644 0.287 0.094 0.488 0.529 0.080 0.067 0.010 0.012 0.047 0.030 0.271
LLAMA3.2 3B 0.842 0.622 0.400 0.733 0.917 0.063 0.052 0.021 0.035 0.042 0.036 0.353
Phi3.5 3B 0.723 0.340 0.188 0.562 0.004 0.0 0.0 0.0 0.0 0.0 0.0 0.002
Fine-Tuned NexusRaven 13B 0.210 0.148 0.082 0.179 N/A N/A 0.160 0.074 0.047 0.124 0.094 0.254
Gorilla v2 7B 0.616 0.005 0.0 0.387 N/A N/A 0.524 0.005 0.0 0.329 0.176 0.518
Table 4: Benchmark Results Table, Fine-tuned model with star* means using Asynchronous Generation with GPT4o calling prediction. Macro Avg. means the arithmetic mean of 3 difficulties.
6.2 Environment
All our local models run with 4-bit Quantization, running on the default Ollama platform settings without any optimization for JSON generation. We do our experiments on NVIDIA-V100 GPU.
6.3 Pipeline
Our evaluation pipeline begins with the creation of prompts based on two templates, as detailed in the appendix. The first template in the appendix B.1 focuses on retrieving the API calling list, which corresponds to the "API" list in the ground truth. This prompt instructs the model to identify which API calls should be used and in what order. The second template is designed to generate the full API calling JSON, including the parameters for each call.
Once the prompts are generated, we run them through each model to obtain predictions. In the first part of the evaluation, where API calls are generated without parameters, we directly calculate the exact match between the predicted API list and the ground truth and make them called API Calling Routing. For the second part, where full JSON outputs are provided, the results are evaluated using the three AST-related scores outlined in Section 5.1: Syntax Validity, Structural Accuracy, and AST Exact Match.
Finally, we employ an LLM-as-a-judge approach, using GPT-4o to calculate a score for comparison, providing a final measure of how well the model’s outputs align with the ground truth. This multi-step process ensures comprehensive evaluation across various levels of output complexity.
6.4 Benchmarks Results
The results presented in Table 6.1 highlight the performance of various models across different aspects of API function calling, including API calling routing accuracy, syntax validity, structural accuracy, and API parameter match through AST evaluation. OpenAI’s models, GPT4o and GPT4o mini, consistently outperform the others, particularly in syntax validity (0.993 and 0.994, respectively) and overall GPT score (0.913 and 0.908). Both models also demonstrate strong structural accuracy and API parameter AST match, especially in easier tasks. And Gemini 1.5 Flash follow closely in these metrics.
Among the large general-purpose open LLMs, LLAMA3.1 (70B) performs well in API calling with an exact match score of 0.945 in easy tasks, though its performance drops significantly in harder cases (0.470). It also achieves the second-highest overall GPT score (0.583), largely due to high syntax validity (0.967). However, its structural accuracy and parameter AST match are weaker, with significant drops in harder tasks. But middle-size LLMs show strong potential ability, very close to the larger group performance such as Gemma2(9B) and xLAM(7B).
The other models and fine-tuned models generally struggle across all indicators. For example, CommandR (35B) shows relatively strong performance in medium API calling tasks (0.877) but performs poorly in structural accuracy (0.189) and API parameter AST match (0.134). Similarly, Mistal models show moderate performance, but the smaller models (e.g., Phi3, LLAMA3.2) display particularly low overall GPT scores and poor performance in most tasks.
Our analysis demonstrates that the Pearson correlation between the "GPT Score" and the "All Avg." column in the "Parameter AST Match" section is 0.934, with a p-value of 4.40e-08. This indicates a very strong positive correlation, suggesting that higher GPT scores are closely associated with better average AST match performance. The results in our table closely align with those from the Berkeley Function Calling Leaderboard [^2] https://gorilla.cs.berkeley.edu/leaderboard.html, which assesses LLMs’ performance in API or function-calling tasks. In both evaluations, models like OpenAI’s GPT4o stand out for their high syntax validity and overall accuracy, as reflected in our table where GPT4o scores above 0.9 in both categories. This matches the leaderboard’s top models, which also excel in AST evaluations and execution accuracy. In contrast, fine-tuned models such as FireFunction V2 in our results show weaker performance in API calling accuracy and AST matching, a trend similarly observed with fine-tuned models like Gorilla OpenFunctions on the Berkeley leaderboard, particularly in more complex API scenarios. Both evaluations emphasize the challenges faced by fine-tuned models in handling complex function scenarios or multi-step API calls, highlighting the need for improvement in these areas.
RQ1: Which LLM have the best performance for function calling in a reality scenario? Answer: Based on the results of our benchmark, we can still claim that OpenAI GPT models are the best solution to solve this kind of challenge. But we can see if the test only by calling the API name list, open-source models can have a closer performance to the state-of-the-art.
6.4.1 Stability Test
In our stability experiments, we ran 5 times referring to the previous study Tian et al. (2023), and the stability results are in Table 7 in Appendix. By these metrics results, we obtain a numerical comparison that reflects the stability differences between outputs.
As mentioned in Section 5.3, a higher Election Stability Score indicates greater absolute consistency in the model’s outputs across multiple runs. A high Levenshtein Stability Score means similar between the same input in text generation output. Commercial models also perform better in Table 7 in Appendix.
RQ2: Which way is the best way to evaluate the API function calling ability of LLMs? Answer: The best way to evaluate the API function calling ability of LLMs is still AST evaluation with parameters. However, our Election Stability Score provides additional insights into output stability, revealing differences that traditional metrics may overlook.
7 Zero Shot Improvement
7.1 Calling + Parameters Asynchronous Generation
We observed that most general-purpose LLMs perform better when generating only API names(routing) rather than both names and parameters simultaneously (see Table 6.1). The added complexity of producing detailed parameters alongside API calls can negatively impact overall performance. Additionally, fine-tuned models struggle with long lists of APIs, limiting their effectiveness in scenarios requiring multiple API calls.
To address these challenges, we propose combining the strengths of general-purpose LLMs and fine-tuned models. Specifically, a general LLM selects the relevant APIs based on the input prompt, leveraging its superior understanding in identifying appropriate API calls. These selected APIs are then provided to a fine-tuned/LAM model, which focuses on generating the correct API calls along with the necessary parameters. This sequential process allows the general LLM to efficiently handle API selection, while the fine-tuned model concentrates on accurately producing API calls and parameters within a more manageable context.
As demonstrated in Table 5 (models marked with an asterisk*), this combined approach with GPT-4o routing significantly improves performance. Separating the tasks of API selection and parameter generation enhances the models’ ability to handle complex API calling tasks more effectively.
Table 5: Asynchronous Generation results for LLMs
Models easy medium hard overall
Fine-Tuned Model w/ GPT routing NexusRaven13B 0.657 0.457 0.188 0.551
Gorilla v27B 0.682 0.005 0.000 0.427
xLAM-fc-7B 0.714 0.462 0.188 0.588
General Large Language Models w/ GPT routing Gemma2:27b 0.633 0.617 0.341 0.595
Gemma2 0.723 0.457 0.164 0.589
llama3.1 0.714 0.457 0.164 0.584
mistral-small 0.728 0.436 0.294 0.602
mistral-nemo 0.712 0.308 0.141 0.541
phi3:14b 0.019 0.005 0.011 0.015
command-r 0.633 0.547 0.223 0.563
llama3.2 0.462 0.297 0.082 0.375
nemotron-mini 0.208 0.01 0.000 0.133
LLM w/ itself as as router Gemma2:27b 0.598 0.59 0.341 0.566
Mistral-small 0.684 0.382 0.235 0.554
Command-r 0.621 0.505 0.247 0.547
7.2 Backward Inference Thinking
To optimize the API selection and calling process, we implemented an approach inspired by CauseJudgerHe and Lu (2024) and Reverse ChainZhang et al. (2023) on the given task. The model is then guided to confirm whether this final API satisfies all the requirements in the question and whether all necessary information is included in the parameters. If the model identifies missing information, it is prompted to think through which additional API calls could provide the missing data. This iterative process continues until all information required by the question is gathered and all requirements are met. By breaking the task into these smaller, more manageable steps, the LLM can effectively plan and execute a series of API calls that better address the complexity of the scenario. This method has led to a performance improvement of approximately 30% in hard-level questions API callings in Table 6 test with GPT-4o/mini, where the challenges of multiple selections and nested calls are more prevalent. This optimization enhances the model’s ability to handle intricate, real-world API calling tasks more effectively.
Table 6: Backward Thinking performance in High difficulty calling in GPT-4o and GPT-4o-mini.
Model API Calling Routing API Calling with Parameters
Original
Backward
Thinking
Original
Backward
Thinking
GPT4o 0.611 0.894 0.388 0.729
GPT4o mini 0.564 0.847 0.364 0.482
RQ3:How to enhance API function calling routing ability for zero/few-shot LLM? Answer: We have tested 2 different ways to enhance the performance of function calling in our test, both can significantly increase performance in function calling routing and JSON generation.
8 Discussion and Conclusion
Our dataset introduces significant challenges, particularly in medium and hard questions, where models must accurately select APIs from a large candidate pool and generate parameters in multi-step and nested contexts. This complexity, absent in simpler datasets, reveals the limitations of fine-tuned models traditionally trained on smaller API sets. The higher variability and complexity of our data underscore the need for more diverse and robust training paradigms.
In evaluating models on API calling routing and parameter JSON generation, we observed distinct behaviors. GPT-4o consistently excelled in both tasks, reinforcing its utility for complex language tasks. Open-source models like LLaMA 3.1 and Gemma2 performed well in API routing but struggled with parameter generation, making them cost-effective options for simpler routing tasks. However, smaller models (<10B parameters) exhibited instability in complex tasks, often producing inconsistent or incomplete outputs, highlighting their limitations in high-accuracy, multi-step API scenarios.
Despite these advancements, none of the models, including GPT-4o, achieved perfect results in handling the hardest scenarios, indicating that fully mastering intricate API calling tasks remains an open challenge.
Conclusion
This work addresses the challenges of API function calling and parameter generation in LLMs by introducing a novel dataset of over 500 APIs and 700 questions, alongside a comprehensive benchmark for general-purpose and fine-tuned models. We also use two optimization strategies asynchronous generation and backward inference to enhance multi-step API selection and parameter generation. Our findings reveal current limitations and open avenues for advancing LLM-driven function calling in theory and practice.
Limitations
While our work presents important advancements in API function calling routing and parameters generation for LLMs, there are several limitations to acknowledge. First, the scenarios we used in our dataset are simplified versions of real-world situations and may not fully capture the complexities that arise in actual applications. This could limit the practical applicability of our results in highly dynamic or unpredictable environments. Additionally, our team has limited expertise in designing datasets across the many diverse domains included in this study, which may have led to oversights in representing certain domain-specific challenges accurately.
Another limitation is related to the performance of the models we tested. While we selected well-regarded LLMs for benchmarking, they may not have performed at their best due to factors such as suboptimal configurations or fine-tuning for different tasks. Lastly, we acknowledge the absence of a mock server in our experimental setup, which could have enhanced the realism of the API calling process. Incorporating a mock server to simulate real-time API interactions would provide a more accurate assessment of how these models handle real-world API calling tasks. Future iterations of this work should address these limitations to provide more comprehensive and applicable insights.
Ethic Statment
In the development of our dataset, we ensured that no private or sensitive information was used, and all data was mock information, generated in compliance with GDPR. This guarantees that no data leakage or privacy concerns are associated with our work. Additionally, we made efforts to minimize the environmental impact of our research by optimizing model training and evaluation processes, thus reducing the carbon footprint associated with our benchmarking tasks. Where possible, we utilized energy-efficient infrastructure to further support this initiative. Our commitment to ethical practices extends across all aspects of this research, ensuring data privacy and environmental responsibility.
Appendix A Meta Data
The first part is the API name, description, and parameters.
"name": "getAccountBalance",
"parameters": ["accountID"],
"description": "Retrieves the current balance for a specific account.",
"returnParameter": {
"Balance": "number"
We then format each question as shown in the example below, which includes the user query, the ground truth API call in JSON format, and the difficulty level:
"id": "ban01",
"question": [
{"role": "user",
"content": "What is the balance for the account with ID 987654?"}],
"ground_truth": {
"API": ["getAccountBalance"],
"parameters":
{"accountID": "987654"}},
"difficulty": "easy"
Appendix B Prompts
B.1 Model prediction
Prompt: Give the API list with descriptions below, then give the question in a chatbot, please give me the correct API that should be called. =======API list start======= {JSON API list} =======API list end======= =======Question start======= {Question} =======Question end======= Given the user question, and the APIs, classify and give a correct API name to call. The answer should be formatted in only one line, and only API names in brackets look like "[’getCustomerDetails’, ’depositFunds’]" or "[’getCustomerDetails’]". NO explanation and NO parameters in the answer!"
Prompt: Give the API list with descriptions below, then give the question in a chatbot, please give me the correct API that should be called. =======API list start======= {JSON API list} =======API list end======= =======Question start======= {Questions} =======Question end======= Given the user question, and the APIs, classify and give a correct API name and parameters to call. The answer should be formatted including API names and parameters in JSON style, which looks like: {’API’: [’getCustomerDetails’, ’depositFunds’], ’parameters’:[{"parameter1ForCall1": "***" },{"parameter1ForCall2": "***", "parameter2ForCall2": "***"}]} If we cannot get some parameter information from the question, set these parameters to "\$\$\$". NO explanation/notes in the answer!
The first prompt is designed for API calling list generating only in the experiment ’API Calling Exact Match’ from Table 6.1. The second prompt is related to generating full JSON output like ground truth in Section A.
B.2 LLM-as-judge
We use the following prompt to generate GPT Score of evaluation in Section 5.2
Prompt: You will be given a predicted API calling JSON and ground truth couple. Your task is to provide a ’total rating’ scoring on how similar the predicted API calling JSON is to ground truth JSON. Give your answer as a float on a scale of 0 to 1 with 3 decimals, where 0 means that the predicted JSON is not related to ground truth, and 1.0 means that the answer is completely the same as ground truth. You should consider if the API calling list is the same or not, and whether all parameters match or not. Only give a score in [[]], looks like [[1.00]]: ======Begin predict JSON====== {predict JSON} ======End of predict JSON====== =====Begin of ground truth JSON===== {ground truth JSON} =====End of ground truth JSON===== Response:
B.3 Backward Thinking
Prompt: Give the API list of a chatbot system with descriptions below, then give the question in chatbot, please first give me the one API that can answer the final task in the question. =======API list start======= {JSON API list} =======API list end======= =======Question start======= {Questions} =======Question end======= Given the user question, and the APIs, classify and give a correct last one task API name and parameters to call. The answer should be formatted into the following JSON format:{format JSON} We will solve the question in the next steps. NO explanation in the answer! ”’
After initial the session, give the following prompt to get a full API calling JSON.
Prompt: Then please based on the answer before, try to find APIs to fill in the missing information needed to complete the final API call list from the previous answer. Please also check if the API list has solved all sub-questions in the question. If not, please add other APIs and parameters related to the question in the correct order. The answer should be formatted in the JSON format used in the previous talk. Check if we can get all input parameters from the questions or not, if we can’t please don’t put <!stop> after the API list. If you think all the information is completed to solve the question, please write a <!stop> after the full API list. NO explanation in the answer! Only give API calling JSON!
Appendix C Case Study
C.1 Hallucination
In one of our test cases, the Phi3:14b model produced an incorrect API function call in response to a baggage tracking scenario. The predicted output was as follows:
"API": ["getLostBaggageReport", "updateBaggageStatus"],
"parameters": [{"baggageId": "BAG123"}, {}]
However, the ground truth was:
"API": ["getBaggageStatus"],
"parameters": [{"baggageId": "BAG123"}]
In this case, the model hallucinated two API calls, "getLostBaggageReport" and "updateBaggageStatus", which were not part of the provided API list. This hallucination led the model to predict incorrect API calls, deviating from the expected function "getBaggageStatus". Although the model correctly captured the parameter baggageId: "BAG123", it introduced an unnecessary second parameter block as an empty dictionary, further reducing the accuracy of the output.
This example highlights a common issue with current large language models in complex tasks: their tendency to hallucinate irrelevant API calls when uncertain. Such behavior emphasizes the need for improved mechanisms to ensure more accurate API function routing and parameter generation in these models.
C.2 JSON Generation
In another example, the Mistral-Nemo model generated an incorrect output, which included unwanted notes in the result, rendering it invalid as a JSON. The predicted output was:
"API": ["getCustomerCreditCards"],
"parameters": [{"customerID": "123456"}]
(Assuming that ATM cards are considered credit cards for this specific API)
The ground truth, however, was:
"API": ["getATMCardList"],
"parameters": [{"accountID": "123456"}]
In this case, the model incorrectly generated an API call for "getCustomerCreditCards" instead of the correct API "getATMCardList". Additionally, the model included an unwanted note—"(Assuming that ATM cards are considered credit cards for this specific API)"—which made the output non-compliant with JSON formatting, as this additional text was outside the structure of the JSON object.
This example illustrates the challenge of maintaining output fidelity in models when they generate explanations or assumptions within the response, which should be avoided in strict JSON-formatted outputs. Such behavior disrupts the automation of API calls and highlights the need for better prompt engineering to ensure models only return valid JSON results without extraneous content.
Appendix D Stability Research
The stability score is calculated in Section 5.3. To give a clear example of calculating the stability of the model’s outputs, we analyze the frequency distribution of the results obtained from 5 times runs. Let’s review the variable settings:
  • $N$ be the total number of outputs (samples).
  • $F_{1}$ be the maximum frequency of any unique output (the most frequent output).
  • $F_{2}$ be the second maximum frequency (the frequency of the second most frequent output).
Explanation:
  • Tie Situations ($F_{1}=F_{2}$): When the maximum frequency is equal to the second maximum frequency, it indicates a tie for the most frequent output. In such cases, the stability score is set to 0 to reflect the lack of a clear majority or consensus.
  • No Tie Situations ($F_{1}>F_{2}$): The numerator $F_{1}-F_{2}$ measures the dominance of the most frequent output over the second most frequent one. The denominator $N-F_{2}$ normalizes this difference relative to the total number of outputs excluding those of the second most frequent output. The resulting score ranges between 0 and 1, where higher values indicate greater stability.
Examples:
  • All Outputs Identical:
    • Results: All outputs are the same (e.g., $[^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime}]$).
    • $F_{1}=N$, $F_{2}=0$ (since there’s only one unique output).
    • Stability Score:
      Stability $\text{Stability Score}=\frac{N-0}{N-0}=\frac{N}{N}=1$
      Indicates perfect stability.
  • Tie Situation (e.g., 2 vs 2 vs 1):
    • Results: Two outputs occur twice, and one occurs once (e.g., $[^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}B^{\prime},^{\prime}C^{\prime}]$).
    • $F_{1}=2$, $F_{2}=2$ (tie between ’A’ and ’B’).
    • Stability Score:
[…]
      Reflects the lack of consensus due to the tie.
  • Minority Advantage (e.g., 2 vs 1 vs 1 vs 1):
    • Results: One output occurs two times, another occurs once each (e.g., $[^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}C^{\prime},^{\prime}D^{\prime}]$).
    • $F_{1}=2$, $F_{2}=1$.
    • Stability Score:
      Stability $\text{Stability Score}=\frac{2-1}{5-1}=\frac{1}{4}\approx 0.25$
      Indicates a little stability.
  • Partial Agreement(Strong Opposition) (e.g., 3 vs 2):
    • Results: One output occurs three times, another occurs twice (e.g., $[^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}B^{\prime}]$).
    • $F_{1}=3$, $F_{2}=2$.
    • Stability Score:
      Stability $\text{Stability Score}=\frac{3-2}{5-2}=\frac{1}{3}\approx 0.333$
      Indicates moderate stability.
  • Partial Agreement(Weak Opposition) (e.g., 3 vs 1 vs 1):
    • Results: One output occurs three times, another occurs once each (e.g., $[^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}C^{\prime}]$).
    • $F_{1}=3$, $F_{2}=1$.
    • Stability Score:
      Stability $\text{Stability Score}=\frac{3-1}{5-1}=\frac{1}{2}\approx 0.5$
      Indicates higher moderate stability.
  • High Majority (e.g., 4 vs 1):
    • Results: One output occurs four times, another occurs once (e.g., $[^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}A^{\prime},^{\prime}B^{\prime}]$).
    • $F_{1}=4$, $F_{2}=1$.
    • Stability Score:
      Stability $\text{Stability Score}=\frac{4-1}{5-1}=\frac{3}{4}=0.75$
      Indicates strong stability.
  • All Outputs Unique:
    • Results: All outputs are different (e.g., $[^{\prime}A^{\prime},^{\prime}B^{\prime},^{\prime}C^{\prime},^{\prime}D^{\prime},^{\prime}E^{\prime}]$).
    • $F_{1}=F_{2}=1$.
    • Stability Score:
[…]
      Reflects complete instability due to lack of any consensus.
Interpretation:
  • Stability Score of 1: Perfect stability; all outputs are identical.
  • Stability Score of 0: No stability; either all outputs are unique or there’s a tie for the most frequent output.
  • Stability Scores Between 0 and 1: Partial stability; higher scores indicate greater agreement among outputs.
Usage:
This stability score provides a quantitative measure of the consistency of the model’s outputs across multiple runs. It accounts for both the dominance of the most frequent output and the impact of significant minority outputs, offering a nuanced assessment of model stability.
Table 7: Stability Test Results
Model Size
Election
Stability Score
Levenshtein
Stability Score
GPT4o N/A 0.674 0.972
GPT4o mini N/A 0.855 0.984
Gemini 1.5 Flash N/A 0.825 0.946
LLAMA3.1 70B 0.407 0.841
LLAMA3.1 8B 0.332 0.740
Mistral-Small 22B 0.208 0.719
Mistral-Nemo 12B 0.365 0.734
CommandR 35B 0.325 0.754
Gemma2 27B 0.609 0.890
Gemma2 9B 0.355 0.864
nemotron-mini 4B 0.013 0.527
LLAMA3.2 3B 0.085 0.613
Phi3.5 3B 0.909 0.637
xLAM-fc 7B 0.782 0.948
Appendix E Example Analysis of Hard and Medium Questions
This section provides an analysis of examples from the CallNavi dataset, focusing on the complexities of medium and hard questions in the bank scenario. The examples illustrate how the tasks test the capabilities of large language models (LLMs) in handling sequential API calls and parameter dependencies.
E.1 Hard Question Example
Question: "I want to travel to Japan, how many Japanese Yen I can spend with my credit card? My customer ID is 123155."
This question involves reasoning across multiple steps and APIs:
  1. 1.
    The model must first retrieve the customer’s credit cards using the getCustomerCreditCards API.
  2. 2.
    From the retrieved cards, the model must fetch details about a specific credit card using the getCreditCardDetails API.
  3. 3.
    Finally, the model must compute the currency exchange rate using the getCurrencyExchangeRates API to determine the spending amount in Japanese Yen.
This task tests the model’s ability to maintain dependencies between APIs, as the output of one API (e.g., credit card number) serves as an input for subsequent calls.
Listing 1: Hard Question Example from the Bank Scenario.
"id": "ban081",
"question": [
"role": "user",
"content": "I want to travel to Japan, how many Japanese Yen I can spend with my credit card? my customer ID is 123155."
"ground_truth": {
"API": ["getCustomerCreditCards", "getCreditCardDetails", "getCurrencyExchangeRates"],
"parameters": [
{ "customerID": "123155" },
{ "creditCardNumber": "\$\$\$" },
{ "currencyPair": "\$\$\$" }
"difficulty": "hard"
E.2 Medium Question Example
Question: "Retrieve details of wire transfer WT987654 and cancel it."
This question involves two sequential API calls:
  1. 1.
    The model must retrieve the details of the wire transfer using the getWireTransferDetails API.
  2. 2.
    The same transfer ID must then be used to cancel the transaction with the cancelWireTransfer API.
This task evaluates the model’s ability to handle basic sequential tasks with consistent parameter reuse across multiple APIs.
Listing 2: Medium Question Example from the Bank Scenario.
"id": "ban069",
"question": [
"role": "user",
"content": "Retrieve details of wire transfer WT987654 and cancel it."
"ground_truth": {
"API": ["getWireTransferDetails", "cancelWireTransfer"],
"parameters": [
{"transferID": "WT987654"},
{"transferID": "WT987654"}
"difficulty": "medium"
E.3 Analysis of Model Challenges
  • Hard Question: The main challenge lies in maintaining dependencies across APIs and correctly passing intermediate results (e.g., credit card number and currency pair).
  • Medium Question: While simpler than the hard question, this task requires consistency in parameter reuse (e.g., the same transferID must be used across both API calls).
These examples illustrate the varying levels of complexity in the CallNavi dataset, testing the models’ abilities to reason through multi-step tasks and maintain logical consistency in API usage.
E.4 Impact of JSON and YAML on Model Performance
To analyze the influence of input and output formats on model performance, we conducted experiments using JSON and YAML, two widely used structured data formats. These formats differ significantly in syntax and structure, which could affect the ability of models to interpret, process, and generate outputs accurately. We tested four configurations:
  • YAML to YAML: Both input and output data are formatted in YAML.
  • JSON to JSON: Both input and output data are formatted in JSON.
  • YAML to JSON: Input data is formatted in YAML, and output data is in JSON.
  • JSON to YAML: Input data is formatted in JSON, and output data is in YAML.
The results of these experiments are shown in Table 8. Models were evaluated on multiple metrics:
  • Syntax Accuracy: Measures the correctness of the generated syntax.
  • Structure Accuracy: Evaluates whether the hierarchical structure of the output matches the ground truth.
  • Task Performance: Evaluates model performance across easy, medium, and hard tasks.
  • Overall Accuracy: Aggregates performance across all task levels.
  • GPT Score: A qualitative metric derived from GPT-based evaluation.
Table 8: Compare JSON or YAML as input/output performance differences.
Input/
Output
Model Syntax Acc Structure Acc easy medium hard overall gpt
YAML to YAML LLAMA3.1 0.525 0.076 0.081 0.042 0.011 0.063 0.183
mistral-small 0 0.183 0.23 0.047 0.035 0.16 0.241
Gemma2:27b 0.938 0.097 0.12 0.037 0 0.085 0.238
command-r 0.883 0.096 0.105 0.042 0.011 0.078 0.241
JSON to JSON LLAMA3.1 0.925 0.207 0.223 0.058 0.059 0.162 0.422
mistral-small 0.986 0.196 0.201 0.106 0.059 0.16 0.417
Gemma2:27b 0.982 0.226 0.217 0.143 0.07 0.181 0.476
command-r 0.969 0.189 0.167 0.095 0.047 0.134 0.4
YAML to JSON LLAMA3.1 0.88 0.194 0.208 0.106 0.023 0.16 0.347
mistral-small 0.995 0.179 0.173 0.112 0.058 0.144 0.333
Gemma2:27b 0.984 0.212 0.195 0.159 0.082 0.172 0.414
command-r 0.967 0.198 0.168 0.122 0.071 0.145 0.322
JSON to YAML LLAMA3.1 0.598 0.104 0.14 0.026 0 0.094 0.32
mistral-small 0 0.218 0.263 0.079 0 0.185 0.332
Gemma2:27b 0.931 0.128 0.153 0.026 0 0.103 0.419
command-r 0.853 0.091 0.123 0.011 0 0.079 0.363
Analysis of Results
1. JSON Outperforms YAML. Across all configurations, models achieved higher syntax, structure, and task-specific accuracies with JSON as both the input and output format. For example, the JSON to JSON configuration resulted in the highest Syntax Accuracy (e.g., 0.986 for Mistral-Small) and Structure Accuracy (e.g., 0.226 for Gemma2:27B), highlighting JSON’s straightforward syntax and reduced ambiguity.
2. YAML Challenges. Models struggled significantly with YAML, particularly in the YAML to YAML configuration, which had the lowest performance across metrics. For instance, LLAMA 3.1 achieved a Syntax Accuracy of 0.525, and Structure Accuracy remained poor across models. YAML’s indentation-sensitive syntax and verbosity likely contribute to these challenges.
3. Mixed Configurations Mitigate Errors. Configurations with mixed input and output formats (e.g., YAML to JSON) performed better than pure YAML setups. JSON as an output format simplified generation tasks, as evidenced by improved metrics compared to YAML outputs.
4. JSON to YAML is Challenging. The JSON to YAML configuration showed decreased performance compared to JSON to JSON, particularly in Syntax Accuracy (e.g., 0.598 for LLAMA 3.1). This indicates that YAML’s complexity as an output format negatively affects model performance, even when the input format is simpler.
Key Observations
  • JSON is more robust for both input and output, offering higher overall performance across tasks.
  • YAML introduces significant challenges in both generation and interpretation, particularly in harder tasks.
  • Mixed configurations, particularly YAML to JSON, mitigate some YAML-related difficulties, highlighting JSON’s simplicity as an output format.
These results underscore the importance of selecting appropriate data formats in real-world applications, as structured formats like YAML can introduce unexpected complexities that affect performance. This study also highlights the potential for optimizing LLMs to better handle YAML and similarly complex formats.
Appendix F Error Analysis
To better understand the limitations of current models in API function-calling tasks, we performed an error analysis on the predictions from our benchmark. This section highlights three typical types of errors observed in the generated outputs: hallucination, unexpected formatting, and logical inconsistencies.
F.1 Hallucination
Hallucination occurs when the model generates API names that do not exist in the provided API candidate pool. This issue often arises in scenarios where the model misinterprets the task requirements or over-generalizes from similar examples. For instance:
  • Example: The model generates an API call to getExchangeRateDetails, whereas the correct API name is getCurrencyExchangeRates.
  • Impact: Such errors result in invalid function calls and demonstrate the model’s difficulty in grounding its outputs to the provided API list.
F.2 Unexpected Formatting
Unexpected formatting errors include outputs containing invalid elements, such as triple backticks (“‘‘‘‘“‘), redundant notes, or commentary text. These errors are common when models generate outputs designed to mimic human explanations but fail to adhere strictly to the expected JSON or YAML format. Examples include:
  • Example: The model outputs:
    ‘‘‘json
    "API": "getCustomerDetails",
    "parameters": { "customerID": "12345"
  • Impact: These formatting errors make the outputs syntactically incorrect and require post-processing to remove extraneous elements.
F.3 Logical Errors in Hard Questions
Logical errors are particularly prevalent in hard questions, where the task involves multiple dependent API calls or complex reasoning. These errors include incorrect sequencing of API calls, failure to propagate parameters correctly, or omitting necessary steps. For example:
  • Example: When asked to retrieve a user’s transaction history and compute their monthly spending, the model retrieves the transactions but fails to invoke an API for computation, leaving the task incomplete.
  • Impact: Logical errors highlight the limitations of current models in handling multi-step reasoning and task dependencies.
F.4 Summary of Error Types
The observed errors reveal challenges in three key areas:
  • Grounding to provided resources (hallucination).
  • Adherence to strict output formats (unexpected formatting).
  • Complex reasoning and multi-step tasks (logical errors).
Addressing these issues will require improving model training data, refining evaluation metrics, and incorporating techniques like grounding and reasoning-specific optimization.